### v2

@@ 3.2 The Suboptimal set partitioning strategy based on Grassmann distance
(Grassmann Distance)Edelman et al. (1998) For two subspaces $U$ and $V$, the Grassmann distance $d_{G}(U,V)$ can be calculated by performing singular value decomposition on the bases of these two subspaces, i.e.,
@@
Based on the indicator function column vector $[I_{j,\ell}]_{\ell=1:n}$ of all sensors $j\in\mathcal{N}_{i}$ in equation (9), the Grassmann distance between the observable spaces of each pair of sensors can be obtained. Then, the problem of assigning completely correlated sensors into the same subset can be described as Problem 3:
@@
$\min_{\theta_{j_{g_{i},1},j_{g_{i},2}}}\ \ \sum_{g_{i}=1:m_{i}}\|\boldsymbol{\theta_{g_{i}}}\|_{2}=\|[\theta_{j_{g_{i},1},j_{g_{i},2}}]_{j_{g_{i},1},j_{g_{i},2}\in\mathcal{N}_{i,g_{i}}}\|_{2}.$ (13)
@@
The Grassmann distance between the observable spaces of sensors $j_{1}$ and $j_{2}$ can be determined using equation (15):
@@
It should be noted that equation (9) has already converted rows containing nonzero elements into binary values, either $1$ or 00. As a result, the Grassmann distance calculated between the indicator function column vectors only has two possible outcomes: either the vectors are perfectly aligned ($\theta_{j_{1},j_{2}}=0$) or they are orthogonal ($\theta_{j_{1},j_{2}}=\pi/2$). Ideally, for any $g_{i}$-th subset, the result of equation (13) should be zero.
@@
8:     Obtain the indicator function column vector $[I_{j,\ell}(k)]_{\ell=1:n}\in\mathbb{R}^{n}$.
@@
Inspired by Lemma 1, we first partition all sensors with the same dimension $Q_{j,o}$ into the same initial set. Then, we only need to apply Algorithm 3.1 separately in each set to partition the subsets. Theoretically, this strategy greatly reduces the computational cost required to calculate the Grassmann distance. The detailed steps are given in Algorithm 3.2.
@@
3:Initialize an empty Map object $rankGroups$.
@@
For the neighbour set $\mathcal{N}_{i}$ of sensor $i$, the set $\mathcal{N}_{i}$ is divided by Algorithm 3.2. At the $l-1$-th selection, suppose that an element $j_{s}^{(l-1)}$ is selected from one of the subset, and only the gains of the remaining sensors in the subset where the element is located are updated. Then, at the $l$-th selection, the distribution ratio vector of the sensor selection of any $g_{i}$-th subset $\mathcal{N}_{i,k,g_{i}}$ is accurate, or the error is tolerable.
@@
For the case where the $l-1$-th selected sensor is completely uncorrelated with the $l$-th selected sensor, that is, there is no intersection in the observation spaces of the two sensors, the gain does not need to be updated at this time, and the distribution ratio vector error is 00.
For the case where the $l-1$-th selected sensor is completely correlated with the $l$-th selected sensor, that is, the observation spaces of the two sensors are exactly the same, the sensor gains are updated. Therefore, the distribution ratio vector error is also 00.
However, for the case where the $l-1$-th selected sensor is partially correlated with the $l$-th selected sensor, that is, the observation spaces of the two sensors intersect but are not exactly the same, so the gains of the two sensors have mutual influence. However, to reduce the computational cost, after the $l-1$th selection, the gain of the $l$th selected sensor is not updated, so there must be an error in the distribution ratio vector. The following proves that the distribution ratio vector error is bounded in this case.
For the $l-1$-th selected sensor $j_{s}^{(l-1)}$ and the $l$-th selected sensor $j_{s}^{(l)}$, the indicator function vectors are partially correlated. Considering the effect of diminishing marginal returns, the gain is most affected when the sensor is initially selected, and gradually decreases for subsequent selections. Therefore, the distribution ratio vector error is tolerable as long as the effect of the $l-1=1$-th selected sensor $j_{s}^{(1)}$ on the distribution ratio vector error of the $l=2$-th selected sensor $j_{s}^{(2)}$ is bounded.
@@
$|p^{\prime}(j_{s}^{(2)})-p^{\prime\prime}(j_{s}^{(2)})|$ (20)
$=$ $\frac{|f_{k}(\{j_{s}^{(2)}\}\cdot\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}\Delta_{j_{s}^{(1)},j^{(2)}}-\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\})\cdot\Delta_{j_{s}^{(1)},j_{s}^{(2)}}|}{(\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\}))\cdot(\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}(f_{k}(\{j^{(2)}\})+\Delta_{j_{s}^{(1)},j^{(2)}}))}$
$\leq$ $\frac{\max_{j^{(2)}}\Delta_{j_{s}^{(1)},j^{(2)}}\cdot|f_{k}(\{j_{s}^{(2)}\}\cdot|\mathcal{N}_{i,k,g_{i}}|-\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\})|}{(\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\}))\cdot(\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}(f_{k}(\{j^{(2)}\})+\Delta_{j_{s}^{(1)},j^{(2)}}))},$
where the inequality holds because of $\Delta_{js^{(1)},js^{(2)}}\leq\max_{j^{(2)}}\Delta_{j_{s}^{(1)},j^{(2)}}=\Delta_{j_{s}^{(1)},j_{max}^{(2)}}$.
First, consider the special case where there is no attack. At this time, the gain of each sensor is only affected by noise. Therefore, from the perspective of the entire time period, the expectation of the gain $f_{k}(\{j^{(2)}\})$ of each element is theoretically 00, that is, $|f_{k}(\{j_{s}^{(2)}\}\cdot|\mathcal{N}_{i,k,g_{i}}|-\sum_{j^{(2)}\in\mathcal{N}_{i,k,g_{i}}}f_{k}(\{j^{(2)}\})|$ is also close to 00, which means that the increase of element $j_{s}^{(1)}$ leads to a negligible change in the gain of element $j^{(2)}$.
@@
For the former, define a $0-1$ binary parameter $\rho$, whose values 00 and $1$ represent the states of sensor $j_{s}^{(2)}$ is under attack (denoted as $j_{s}^{(2),u}$) and without attack (denoted as $j_{s}^{(2),w}$), respectively. At this time, the following derivation is obtained
@@
$=$ $|f_{k}(\{j_{s}^{(2)}\}\cdot|\mathcal{N}_{i,k,g_{i}}|-\sum_{j^{(2),w}\in\mathcal{N}_{i,k,g_{i}}\backslash I_{i,k,g_{i}}}f_{k}(\{j^{(2),w}\})-\sum_{j^{(2),u}\in I_{i,k,g_{i}}}f_{k}(\{j^{(2),u}\})|$
$\leq$ $\rho\cdot\left(|\mathcal{N}_{i,k,g_{i}}|-q_{i,g_{i}}\right)|\min_{j^{(2),w}\in\mathcal{N}_{i,k,g_{i}}\backslash I_{i,k,g_{i}}}f_{k}(\{j^{(2),w}\})-\max_{j^{(2),u}\in I_{i,k,g_{i}}}f_{k}(\{j^{(2),u}\})|$
@@
where $I_{i,k,g_{i}}$ in the first equation represents the set of all attacked sensors in the set $N_{i,k,g_{i}}$, and the first inequality holds because the maximum error that the selected sensor may bring is the maximum gain in the attacked sensor set and the minimum gain in the normal sensor set. In the second equation, $j_{max}^{(2),u}$ and $j_{min}^{(2),w}$ are equal to $\max_{j^{(2),u}\in I_{i,k,g_{i}}}f_{k}(\{j^{(2),u}\})$ and $\min_{j^{(2),w}\in\mathcal{N}_{i,k,g_{i}}\backslash I_{i,k,g_{i}}}f_{k}(\{j^{(2),w}\})$, respectively.
In addition, the parameter $\mu_{ij}^{(u)}(k)$ ($\mu_{ij}^{(w)}(k)$) is defined as the estimation error when the communication link between sensor $j$ and $i$ is under attack (without attack), which only depends on the network effect and the measurement noise and process noise of sensors $i$ and $j$. By taking the expectation of equation (21) and omitting the subscripts $max$ and $min$, we can get:
@@
$\leq$ $\max\left\{\lfloor\frac{|\mathcal{N}_{i}|-q_{i}}{m_{i}}\rfloor+\mathbb{I}(g_{i}\leq(|\mathcal{N}_{i}|\ mod\ m_{i}|)),\frac{q_{i}}{m_{i}}\right\}\cdot\frac{(\sum_{l=1}^{q_{i}}\phi_{i,q_{i}}(k)/q_{i}+2||\mu_{i}(k)||^{2}_{\infty})}{\bar{\mu_{i}}(k)},$
where the first inequality is obtained by multiplying both the numerator and the denominator by $|f_{k}(\{j^{(2),u}\})+f_{k}(\{j^{(2),w}\})|$. The first equality holds because the values of $f_{k}(\{j^{(2),u}\})$ and $f_{k}(\{j^{(2),w}\})$ are essentially related to the attack signal $z_{i,j^{(2),u}}$ and the estimation error $\mu_{i,j^{(2),u}}^{u}$ ($\mu_{i,j^{(2),w}}^{w}$). The third inequality holds because $z_{i,j^{(2),u}}(k)$ and $\mu_{i,j^{(2),u}}^{u}(k)$ are completely independent over the entire time period, and the expectation of the attack signal $z_{i,j^{(2),u}}^{2}$ on any sensor $j^{(2),u}$ is approximately $1/q_{i}$, which is the sum of the average malicious perturbation power of all attacks, expressed as $\mathbb{E}[z_{i,j^{(2),u}}(k)^{2}]=\sum_{l=1}^{q_{i}}\phi_{i,q_{i}}(k)/q_{i}$. In addition, the estimated error $\mu_{ij}\left(k\right)$ satisfies $||\mu_{i}\left(k\right)||_{\infty}=\max_{j\in\mathcal{N}_{i}}\left\{||\mu_{ij}\left(k\right)||_{\infty}\right\}$ and the estimation error $\mu_{ij}\left(k\right)$ satisfies $|z_{i,j^{(2),u}}(k)|\geq\left\|\mu_{i}(k)\right\|_{\infty}\geq|\sum_{j\in\mathcal{N}_{i}}{\mu_{ij}(k)}|/|\mathcal{N}_{i}|=\bar{\mu_{i}}$.
@@
Considering the two parts of equation (20), ideally, $|p^{\prime}(j_{s}^{(2)})-p^{\prime\prime}(j_{s}^{(2)})|\rightarrow 0$, because a bounded positive number multiplied by an infinitesimal number tends to 00. In practice, we only need to ensure $|p^{\prime}(j_{s}^{(2)})-p^{\prime\prime}(j_{s}^{(2)})|\leq\epsilon$, which is obviously achievable. This also means that based on the proposed set partitioning strategy, only the gains of the remaining sensors in the subset where the selected sensor is located need to be updated, instead of the gains of all the remaining sensors, which will significantly reduce the calculation of the gains. This complete the proof.

### v3

@@ 3.2 The set partitioning strategy based on Grassmann distance
(Grassmann Distance) (Edelman et al., 1998) For two subspaces $U$ and $V$, the Grassmann distance $d_{G}(U,\,V)$ can be calculated by performing singular value decomposition on the bases of these two subspaces, i.e.,
@@
Based on the indicator function column vector $I_{j,\ \ell}]_{\ell=1:n}$ of all sensors $j\in\mathcal{N}_{i}$ in equation (9), the Grassmann distance between the observable spaces of each pair of sensors can be obtained. Then, the problem of assigning completely correlated sensors into the same subset can be described as Problem 3:
@@
$\min_{\theta_{j_{g_{i},\,1},\,j_{g_{i},\,2}}}\,\ \sum_{g_{i}=1:m_{i}}\|[\theta_{j_{g_{i},\,1},\,j_{g_{i},\,2}}]_{j_{g_{i},\,1},j_{g_{i},\,2}\in\mathcal{N}_{i,\,g_{i}}}\|_{2}.$ (13)
@@
The Grassmann distance between the observable spaces of sensors $j_{1}$ and $j_{2}$ can be determined by
@@
It should be noted that equation (9) has already converted rows containing nonzero elements into binary values, either $1$ or 0. As a result, the Grassmann distance calculated between the indicator function column vectors only has two possible outcomes: either the vectors are perfectly aligned ($\theta_{j_{1},\,j_{2}}=0$) or they are orthogonal ($\theta_{j_{1},\,j_{2}}=\pi/2$). Ideally, for any $g_{i}$-th subset, the result of equation (13) should be zero.
@@
8:  Obtain the indicator function column vector $I_{j,\,\ell}(k)]_{\ell=1:n}\in\mathbb{R}^{n}$.
@@
Inspired by Lemma 1, we first partition all sensors with the same dimension $Q_{j,\,o}$ into the same initial set. Then, we only need to apply Algorithm 3.1 separately in each set to partition the subsets. Theoretically, this strategy greatly reduces the computational cost required to calculate the Grassmann distance. The details are given in Algorithm 3.2.
@@
3:Initialize an empty Map object rankGroupsrankGroups.
@@
On this basis, the ADS algorithm in the literature Suo et al. (2024c) can be deployed on each subset. At this time, the ADS algorithm becomes a distributed case, that is, the D-ADS algorithm. Theoretical results show that the D-ADS algorithm guarantees the same performance lower bound for each subset as the ADS algorithm. The proof process is omitted here. Please refer to the APPENDIX B.
@@ 3.3 Theoretical performance
For the neighbour set $\mathcal{N}_{i}$ of sensor $i$, the set $\mathcal{N}_{i}$ is divided by Algorithm 3.2. At the l1l-1-th selection, suppose that an element $j_{s}^{(l-1)}$ is selected from one of the subset, and only the gains of the remaining sensors in the subset where the element is located are updated. Then, at the $l$-th selection, the distribution ratio vector of the sensor selection of any $g_{i}$-th subset $\mathcal{N}_{i,\,k,\,g_{i}}$ is accurate, or the error is tolerable.
@@
For the case where the l1l-1-th selected sensor is completely uncorrelated with the $l$-th selected sensor, that is, there is no intersection in the observation spaces of the two sensors, the gain does not need to be updated at this time, and the distribution ratio vector error is 0.
For the case where the l1l-1-th selected sensor is completely correlated with the $l$-th selected sensor, that is, the observation spaces of the two sensors are exactly the same, the sensor gains are updated. Therefore, the distribution ratio vector error is also 0.
However, for the case where the l1l-1-th selected sensor is partially correlated with the $l$-th selected sensor, that is, the observation spaces of the two sensors intersect but are not exactly the same, so the gains of the two sensors have mutual influence. However, to reduce the computational cost, after the l1l-1th selection, the gain of the $l$th selected sensor is not updated, so there must be an error in the distribution ratio vector. The following proves that the distribution ratio vector error is bounded in this case.
For the l1l-1-th selected sensor $j_{s}^{(l-1)}$ and the $l$-th selected sensor $j_{s}^{(l)}$, the indicator function vectors are partially correlated. Considering the effect of diminishing marginal returns, the gain is most affected when the sensor is initially selected, and gradually decreases for subsequent selections. Therefore, the distribution ratio vector error is tolerable as long as the effect of the l1=1l-1=1-th selected sensor $j_{s}^{(1)}$ on the distribution ratio vector error of the $l=2$-th selected sensor $j_{s}^{(2)}$ is bounded.
@@
$\bigl|p^{\prime}(j_{s}^{(2)})-p^{\prime\prime}(j_{s}^{(2)})\bigr|$ (20)
$=$ $\frac{\bigl|f_{k}(\{j_{s}^{(2)}\})\cdot\sum_{j^{(2)}\in\mathcal{N}_{i,\,k,\,g_{i}}}\Delta_{\,j_{s}^{(1)},\,j^{(2)}}\;-\;\sum_{j^{(2)}\in\mathcal{N}_{i,\,k,\,g_{i}}}f_{k}(\{j^{(2)}\})\cdot\Delta_{\,j_{s}^{(1)},\,j_{s}^{(2)}}\bigr|}{\bigl(\sum_{j^{(2)}\in\mathcal{N}_{i,\,k,\,g_{i}}}f_{k}(\{j^{(2)}\})\bigr)\,\cdot\,\bigl(\sum_{j^{(2)}\in\mathcal{N}_{i,\,k,\,g_{i}}}(f_{k}(\{j^{(2)}\})+\Delta_{\,j_{s}^{(1)},\,j^{(2)}})\bigr)}$
$\leq$ $\frac{\max_{j^{(2)}}\Delta_{\,j_{s}^{(1)},\,j^{(2)}}\,\cdot\,\bigl|f_{k}(\{j_{s}^{(2)}\})\cdot|\mathcal{N}_{i,\,k,\,g_{i}}|\;-\;\sum_{j^{(2)}\in\mathcal{N}_{i,\,k,\,g_{i}}}f_{k}(\{j^{(2)}\})\bigr|}{\bigl(\sum_{j^{(2)}\in\mathcal{N}_{i,\,k,\,g_{i}}}f_{k}(\{j^{(2)}\})\bigr)\,\cdot\,\bigl(\sum_{j^{(2)}\in\mathcal{N}_{i,\,k,\,g_{i}}}(f_{k}(\{j^{(2)}\})+\Delta_{\,j_{s}^{(1)},\,j^{(2)}})\bigr)},$
where the inequality holds because of $\Delta_{j_{s}^{(1)},\,j_{s}^{(2)}}\leq\max_{j^{(2)}}\Delta_{j_{s}^{(1)},\,j^{(2)}}=\Delta_{j_{s}^{(1)},\,j_{max}^{(2)}}$.
First, consider the special case where there is no attack. At this time, the gain of each sensor is only affected by noise. Therefore, from the perspective of the entire time period, the expectation of the gain $f_{k}(\{j^{(2)}\})$ of each element is theoretically 0, that is, $|f_{k}(\{j_{s}^{(2)}\}\cdot|\mathcal{N}_{i,\,k,\,g_{i}}|-\sum_{j^{(2)}\in\mathcal{N}_{i,\,k,\,g_{i}}}f_{k}(\{j^{(2)}\})|$ is also close to 0, which means that the increase of element $j_{s}^{(1)}$ leads to a negligible change in the gain of element $j^{(2)}$.
@@
For the former, define a 010-1 binary parameter $\rho$, whose values 0 and $1$ represent the states of sensor $j_{s}^{(2)}$ is under attack (denoted as $j_{s}^{(2),\,u}$) and without attack (denoted as $j_{s}^{(2),\,w}$), respectively. At this time, the following derivation is obtained
@@
[…]
@@
where $I_{i,\,k,\,g_{i}}$ in the first equation represents the set of all attacked sensors in the set $N_{i,\,k,\,g_{i}}$, and the first inequality holds because the maximum error that the selected sensor may bring is the maximum gain in the attacked sensor set and the minimum gain in the normal sensor set. In the second equation, $j_{max}^{(2),\,u}$ and $j_{min}^{(2),\,w}$ are equal to $\max_{j^{(2),\,u}\in I_{i,\,k,\,g_{i}}}f_{k}(\{j^{(2),\,u}\})$ and $min_{j^{(2),\,w}\in\mathcal{N}_{i,\,k,\,g_{i}}\backslash I_{i,\,k,\,g_{i}}}f_{k}(\{j^{(2),\,w}\})$, respectively.
In addition, the parameter $\mu_{ij}^{(u)}(k)$ ($\mu_{ij}^{(w)}(k)$) is defined as the estimation error when the communication link between sensor $j$ and $i$ is under attack (without attack), which only depends on the network effect and the measurement noise and process noise of sensors $i$ and $j$. By taking the expectation of equation (21) and omitting the subscripts maxmax and minmin, we can get
@@
[…]
@@
Considering the two parts of equation (20), ideally, $|p^{\prime}(j_{s}^{(2)})-p^{\prime\prime}(j_{s}^{(2)})|\rightarrow 0$, because a bounded positive number multiplied by an infinitesimal number tends to 0. In practice, we only need to ensure $|p^{\prime}(j_{s}^{(2)})-p^{\prime\prime}(j_{s}^{(2)})|\leq\epsilon$, which is obviously achievable. This also means that based on the proposed set partitioning strategy, only the gains of the remaining sensors in the subset where the selected sensor is located need to be updated, instead of the gains of all the remaining sensors, which will significantly reduce the calculation of the gains. This complete the proof.
3.4 The impact of different gain update methods on algorithm performance
In this section, $5$ different gain update methods are compared in terms of computational complexity, distribution ratio vector accuracy, and performance. The considered methods are as follows:
  • (1)
    Method $1$: Divide the total sensor set into $m_{i}$ subsets by Algorithm 3.2, and update the gains of all fully correlated elements in the subset after each sensor selection.
  • (2)
    Method $2$: Divide the total sensor set into $m_{i}$ subsets by Algorithm 3.2, and update the gains of all related elements after each sensor selection.
  • (3)
    Method $3$: Randomly divide the total sensor set into $m_{i}$ subsets, and update the gains of all fully correlated sensors in the subset after each sensor selection.
  • (4)
    Method $4$: Randomly divide the total sensor set into $m_{i}$ subsets, and update the gains of all related sensors in the subset after each sensor selection.
  • (5)
    Method $5$: without divide the totaol sensor set, and update the gains of all related sensors after each sensor selection.
To ensure fairness in the comparison, the following conditions are imposed. For all methods requiring partitioning, the total sensor set is divided into same number of subsets; all estimated error vectors are processed using the observable discrimination matrix; malicious information is selected based on the distribution ratio vector; and the attack intensity across subsets is completely balanced. The comparison results are summarized in Table 1.
Table 1: Effect of different gain update methods on computational complexity, distribution ratio vector accuracy, and performance
number of computational distribution ratio lower bound performance
subsets complexity vector accuracy overall subset
Method 1 $m_{i}$ larger accurate near-optimal 11/e1-1/e
Method 2 $m_{i}$
large
most accurate optimal 11/e1-1/e
Method 3 $m_{i}$ larger random random 11/e1-1/e
Method 4 $m_{i}$
larger
random
random 11/e1-1/e
Method 5 1 largest most accurate optimal, 11/e1-1/e none
Method $5$ essentially corresponds to the ADS algorithm proposed in Suo et al. (2024c). Among the five gain update methods, it exhibits the highest computational complexity, since it involves normalizing high-dimensional vectors and updating the gains of all correlated elements. Consequently, this method achieves the most accurate distribution ratio vector and delivers the best overall performance. In the subsequent discussion, Method $5$ will be used as the benchmark.
Theoretically, methods $2$ and $5$ achieve comparable distribution ratio vector accuracy and overall performance, but method $2$ requires lower computational complexity (e.g., lower-dimensional normalization). The key distinction between methods $1$ and $2$ lies in whether correlated elements in other subsets are updated after each sensor selection; as a result, method $1$ has lower complexity but also reduced accuracy and performance. Similarly, the difference between methods $3$ and $4$ is the number of updated element gains, making method $4$ more complex but more accurate. For methods 1144, subset-level performance is guaranteed by Theorem 2, with a theoretical lower bound of 11/e1-1/e. However, their overall performance remains inferior to method $5$, reflecting the inherent gap between distributed and centralized algorithms. Notably, method $1$ is expected to deliver better and more stable performance than random partitioning.
Remark 4
In fact, building on the proposed set partitioning strategy, the distributed structure is naturally amenable to parallel computing frameworks, as processing within each subset is independent and cross-subset synchronization is minimal, so further efficiency gains can be expected in practical implementations. Meanwhile, even without parallel acceleration, the per-iteration computational cost scales as $O(1/m)$ with the number of subsets $m$, where $m$ denotes the number of subsets.